### v2

@@ III.3 $D^{*}\bar{B}^{*}$ systems
$\mathcal{M}^{(2)}_{c3.5}=$ $-i4\frac{g_{1}^{2}}{f^{4}}(d-3)[A_{1}(q_{0}J^{T}_{22}+J^{T}_{24}+q_{0}J^{T}_{32}+J^{T}_{33})-A_{5}(q_{0}J^{T}_{11}+2q_{0}J^{T}_{22}+J^{T}_{24}+{q}_{0}J^{T}_{32}+J^{T}_{33})]\mathcal{O}^{a}_{1}$
$-i4\frac{g^{2}_{1}}{f^{4}}(d-3)\{A_{1}[({q}_{0}\vec{q}^{2}J^{T}_{22}+\vec{q}^{2}J^{T}_{24}+(2-d)q_{0}J^{T}_{31}+q_{0}\vec{q}^{2}J^{T}_{32}+\vec{q}^{2}J^{T}_{33}+(2-d)q_{0}J^{T}_{34})]$
$-A_{5}[q_{0}\vec{q}^{2}J^{T}_{11}+(2-d)q_{0}J^{T}_{21}+\vec{q}^{2}J^{T}_{24}+(2-d)q_{0}J^{T}_{31}+q_{0}\vec{q}^{2}J^{T}_{32}+\vec{q}^{2}J^{T}_{33}+(2-d)q_{0}J^{T}_{34}]\}\mathcal{O}_{1},$ (91)
$\mathcal{M}^{(2)}_{c3.6}=$ $-4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}[(\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{B}_{41}+\mathcal{O}^{b}_{3}(J^{B}_{21}+2J^{B}_{31}+J^{B}_{42})+(\mathcal{O}^{a}_{1}+\mathcal{O}^{b}_{1}+\mathcal{O}^{a}_{2}+\mathcal{O}^{b}_{2})(J^{B}_{31}+J^{B}_{42})+\mathcal{O}^{a}_{3}J^{B}_{42}$
@@
$\mathcal{M}^{(2)}_{c3.7}=$ $-4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)^{2}A_{1}\{[(d-2)(d-1)\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3}]J^{B}_{41}+({\vec{q}}^{4}\mathcal{O}_{1}+{\vec{q}}^{2}\mathcal{O}^{a}_{1}+{\vec{q}}^{2}\mathcal{O}^{b}_{1}+\mathcal{O}^{c}_{1})(J^{B}_{22}+2J^{B}_{32}+J^{B}_{43})$
$-({\vec{q}}^{2}\mathcal{O}_{1}+\mathcal{O}^{a}_{1}+\mathcal{O}^{b}_{1}-\mathcal{O}^{a}_{3})J^{B}_{21}-[2d{\vec{q}}^{2}\mathcal{O}_{1}+(d+2)\mathcal{O}^{a}_{1}+(d+2)\mathcal{O}^{b}_{1}-\mathcal{O}^{a}_{2}-\mathcal{O}^{b}_{2}-\mathcal{O}^{b}_{3}](J^{B}_{31}+J^{B}_{42})$
$+\mathcal{O}^{b}_{3}J^{B}_{31}+\mathcal{O}^{a}_{3}J^{B}_{42}\},$ (93)
$\mathcal{M}^{(2)}_{c3.8}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{(-d\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{B}_{41}+({\vec{q}}^{2}\mathcal{O}^{b}_{1}+\mathcal{O}^{c}_{3})(J^{B}_{22}+2J^{B}_{32}+J^{B}_{43})+(\mathcal{O}^{b}_{2}-\mathcal{O}^{b}_{1})J^{B}_{21}$
$+[{\vec{q}}^{2}\mathcal{O}_{1}+\mathcal{O}^{a}_{1}-(d+2)\mathcal{O}^{b}_{1}+\mathcal{O}^{b}_{2}+\mathcal{O}^{a}_{3}+\mathcal{O}^{b}_{3}](J^{B}_{31}+J^{B}_{42})+\mathcal{O}^{b}_{2}J^{B}_{31}+\mathcal{O}^{a}_{2}J^{B}_{42}\},$ (94)
$\mathcal{M}^{(2)}_{c3.9}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{(-d\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{B}_{41}+({\vec{q}}^{2}\mathcal{O}^{a}_{1}+\mathcal{O}^{c}_{3})(J^{B}_{22}+2J^{B}_{32}+J^{B}_{43})+(\mathcal{O}^{a}_{2}-\mathcal{O}^{a}_{1})J^{B}_{21}$
$+[{\vec{q}}^{2}\mathcal{O}_{1}-(d+2)\mathcal{O}^{a}_{1}+\mathcal{O}^{b}_{1}+\mathcal{O}^{a}_{2}+\mathcal{O}^{a}_{3}+\mathcal{O}^{b}_{3}](J^{B}_{31}+J^{B}_{42})+\mathcal{O}^{a}_{2}J^{B}_{31}+\mathcal{O}^{b}_{2}J^{B}_{42}\},$ (95)
$\mathcal{M}^{(2)}_{c3.10}=$ $~{}5\mathcal{M}^{(2)}_{c3.6}/.\{J^{B}\to J^{R}\},\quad\mathcal{M}^{(2)}_{c3.11}=5\mathcal{M}^{(2)}_{c3.7}/.\{J^{B}\to J^{R}\},$ (96)
$\mathcal{M}^{(2)}_{c2.12}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{(-d\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{R}_{41}+({\vec{q}}^{2}\mathcal{O}^{b}_{1}+\mathcal{O}^{c}_{3})(J^{R}_{22}+2J^{R}_{32}+J^{R}_{43})+(\mathcal{O}^{b}_{2}-\mathcal{O}^{b}_{1})J^{R}_{21}$
$+[{\vec{q}}^{2}\mathcal{O}_{1}+\mathcal{O}^{a}_{1}-(d+2)\mathcal{O}^{b}_{1}+\mathcal{O}^{a}_{2}+\mathcal{O}^{b}_{2}+\mathcal{O}^{b}_{3}](J^{R}_{31}+J^{R}_{42})+\mathcal{O}^{b}_{3}J^{R}_{31}+\mathcal{O}^{a}_{3}J^{R}_{42}\},$ (97)
$\mathcal{M}^{(2)}_{c3.13}=$ $~{}4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{(-d\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{R}_{41}+({\vec{q}}^{2}\mathcal{O}^{b}_{1}+\mathcal{O}^{c}_{3})(J^{R}_{22}+2J^{R}_{32}+J^{R}_{43})+(\mathcal{O}^{b}_{2}-\mathcal{O}^{b}_{1})J^{R}_{21}$
$+[{\vec{q}}^{2}\mathcal{O}_{1}+\mathcal{O}^{a}_{1}-(d+2)\mathcal{O}^{b}_{1}+\mathcal{O}^{a}_{2}+\mathcal{O}^{b}_{2}+\mathcal{O}^{a}_{3}](J^{R}_{31}+J^{R}_{42})+\mathcal{O}^{a}_{3}J^{R}_{31}+\mathcal{O}^{b}_{3}J^{R}_{42}\},$ (98)
@@
To further evaluate the calculated amplitudes, we should deal with the terms (54) according to the partial wave we considered. In $S$-wave, for the terms like $(q\cdot\epsilon_{i})(q\cdot\epsilon_{j}^{*})(\epsilon_{k}\cdot\epsilon_{l}^{*})$ we can make the following substitutions:
@@
Table 5: The values of the products of polarization vectors in the $S$-wave effective potentials with total angular momentum $J=0$, $1$, and $2$, respectively Wang:2018atz .
@@
     $\mathcal{O}_{1}$      $1$      $1$      $1$
     $\mathcal{O}_{2}$      $1$      $-1$      $1$
     $\mathcal{O}_{3}$      $3$      00      00
Table 6: The coefficients appearing in the contact term amplitudes at NLO (Eqs. (57)-(74)).
@@
$A_{a3.1}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{1}{4}(3D_{b}-E_{b})$ $\frac{1}{4}(3D_{b}-E_{b}$) $\frac{3}{4}(D_{a}+E_{a})$ $\frac{3}{4}(D_{b}+E_{b})$ $\frac{3}{4}(D_{b}+E_{b})$ 00 00
$A_{a3.2}$ $\frac{1}{4}(3D_{a}-E_{a})$ 00 00 $\frac{3}{4}(D_{a}+E_{a})$ 00 00 $\delta_{2}$ $\delta_{2}$
$A_{a3.3(4)}$ $\frac{1}{4}(3D_{b}-E_{b})$ $\frac{1}{4}(3D_{b}-E_{b})$ 00 $\frac{3}{4}(D_{b}+E_{b})$ $\frac{3}{4}(D_{b}+E_{b})$ 00 $0(\delta_{2})$ $\delta_{2}(0)$
$A_{a3.5}$ 00 00 00 00 00 00 00 00
$A_{a3.6}$ 00 00 00 00 00 00 $\delta_{1}$ $\delta_{1}$
$A_{a3.7(8)}$ 00 00 00 00 00 00 $0(\delta_{1})$ $\delta_{1}(0)$
$A_{a3.9}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{1}{4}[(d-3)(D_{b}+E_{b})$ $+(D_{a}+E_{a})]$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\frac{-3}{4}[(d-3)(D_{b}-3E_{b})$ $+(D_{a}-3E_{a})]$ 00 00
$A_{a3.10}$ $\frac{1}{4}(D_{b}+E_{b})$ 00 00 $\frac{-3}{4}(D_{b}-3E_{b})$ 00 00 $\delta_{1}$ $\delta_{2}$
$A_{a3.11(12)}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{b}+E_{b})$ 00 $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 $0(\delta_{1})$ $\delta_{2}(0)$
$A_{a3.13}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{1}{4}[(d-3)(D_{b}+E_{b})$ $+(D_{a}+E_{a})]$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\frac{-3}{4}[(d-3)(D_{b}-3E_{b})$ $+(D_{a}-3E_{a})]$ 00 00
$A_{a3.14}$ $\frac{1}{4}(D_{b}+E_{b})$ 00 00 $\frac{-3}{4}(D_{b}-3E_{b})$ 00 00 $\delta_{1}$ $\delta_{2}$
$A_{a3.15(16)}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{b}+E_{b})$ 00 $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 $0(\delta_{1})$ $\delta_{2}(0)$
$A_{a3.(17+18)}$ $D_{a}+E_{a}$ $D_{b}+E_{b}$ $D_{b}+E_{b}$ $D_{a}-3E_{a}$ $D_{b}-3E_{b}$ $D_{b}-3E_{b}$ 00 $\delta_{2}$
$A_{a3.(19+20)}$ $D_{a}+E_{a}$ $D_{b}+E_{b}$ $D_{b}+E_{b}$ $D_{a}-3E_{a}$ $D_{b}-3E_{b}$ $D_{b}-3E_{b}$ 00 $\delta_{1}$
Table 7: The constants $A$ (as well as $\omega_{1,2}$) appearing in the OPE amplitudes (Eqs. (75)-(86)).
@@
$I=1$ $-\frac{1}{16}$ $-\frac{1}{16}$ $-\frac{1}{16}$ $-\frac{1}{16}$ $-\frac{1}{12}$ $-\frac{1}{12}$ $\frac{1}{4}$ $\frac{1}{4}$ $\frac{1}{4}$ 00 00
$I=0$ $\frac{3}{16}$ $\frac{3}{16}$ $\frac{3}{16}$ $\frac{3}{16}$ $\frac{1}{4}$ $\frac{1}{4}$ $-\frac{3}{4}$ $-\frac{3}{4}$ $-\frac{3}{4}$ 00 00
$\omega_{1}$ 00 $\delta_{2}(0)$ 00 $0(\delta_{1})$ 00 00 00 00 00 00 00
$\omega_{2}$ 00 $0(\delta_{2})$ 00 $\delta_{1}(0)$ 00 00 00 $\delta_{1}$ $\delta_{2}$ 00 00
Table 8: The constants appearing in the TPE amplitudes (Eqs. (87)-(98)). Note that we have $A_{51}=A_{15}$.
@@
       $A_{c3.1}$        $\frac{1}{16}$        $\frac{1}{16}$        $-\frac{1}{16}$        $-\frac{3}{16}$        $-\frac{3}{16}$        $\frac{3}{16}$        00        00
       $A_{c3.2}$        $\frac{i}{8}$        $-\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        $\delta_{2}$        00
       $A_{c3.3}$        $\frac{i}{8}$        $-\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        00        00
       $A_{c3.4}$        $\frac{i}{8}$        $-\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        $\delta_{1}$        00
       $A_{c3.5}$        $\frac{i}{8}$        $-\frac{i}{8}$        00        $-\frac{3i}{8}$        $\frac{3i}{8}$        00        00        00
       $A_{c3.6}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        $\delta_{1}$        $\delta_{2}$
       $A_{c3.7}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        00        00
       $A_{c3.8}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        $\delta_{1}$        00
       $A_{c3.9}$        $\frac{1}{16}$        00        00        $\frac{9}{16}$        00        00        00        $\delta_{2}$
       $A_{c3.10}$        $\frac{5}{16}$        00        00        $-\frac{3}{16}$        00        00        $\delta_{1}$        $\delta_{2}$
       $A_{c3.11}$        $\frac{5}{16}$        00        00        $-\frac{3}{16}$        00        00        00        00
       $A_{c3.12}$        $\frac{5}{16}$        00        00        $-\frac{3}{16}$        00        00        $\delta_{1}$        00
       $A_{c3.13}$        $\frac{5}{16}$        00        00        $-\frac{3}{16}$        00        00        00        $\delta_{2}$

### v3

@@ III.3 $D^{*}\bar{B}^{*}$ systems
$\mathcal{M}^{(2)}_{c3.5}$= $-i4\frac{g_{1}^{2}}{f^{4}}(d-3)[A_{1}(q_{0}J^{T}_{22}+J^{T}_{24}+q_{0}J^{T}_{32}+J^{T}_{33})$
$-A_{5}(q_{0}J^{T}_{11}+2q_{0}J^{T}_{22}+J^{T}_{24}+{q}_{0}J^{T}_{32}+J^{T}_{33})]\mathcal{O}^{a}_{1}$
$-i4\frac{g^{2}_{1}}{f^{4}}(d-3)\{A_{1}[({q}_{0}\vec{q}^{2}J^{T}_{22}+\vec{q}^{2}J^{T}_{24}$
$+(2-d)q_{0}J^{T}_{31}+q_{0}\vec{q}^{2}J^{T}_{32}+\vec{q}^{2}J^{T}_{33}+(2-d)q_{0}J^{T}_{34})]$
$-A_{5}[q_{0}\vec{q}^{2}J^{T}_{11}+(2-d)q_{0}J^{T}_{21}+\vec{q}^{2}J^{T}_{24}$
$+(2-d)q_{0}J^{T}_{31}+q_{0}\vec{q}^{2}J^{T}_{32}+\vec{q}^{2}J^{T}_{33}$
$+(2-d)q_{0}J^{T}_{34}]\}\mathcal{O}_{1},$ (91)
$\mathcal{M}^{(2)}_{c3.6}$= $-4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}A_{1}[(\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{B}_{41}$
$+\mathcal{O}^{b}_{3}(J^{B}_{21}+2J^{B}_{31}+J^{B}_{42})$
$+(\mathcal{O}^{a}_{1}+\mathcal{O}^{b}_{1}+\mathcal{O}^{a}_{2}+\mathcal{O}^{b}_{2})(J^{B}_{31}+J^{B}_{42})+\mathcal{O}^{a}_{3}J^{B}_{42}$
@@
$\mathcal{M}^{(2)}_{c3.7}$= $-4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)^{2}A_{1}\{[(d-2)(d-1)\mathcal{O}_{1}$
$+\mathcal{O}_{2}+\mathcal{O}_{3}]J^{B}_{41}+({\vec{q}}^{4}\mathcal{O}_{1}+{\vec{q}}^{2}\mathcal{O}^{a}_{1}$
$+{\vec{q}}^{2}\mathcal{O}^{b}_{1}+\mathcal{O}^{c}_{1})(J^{B}_{22}+2J^{B}_{32}+J^{B}_{43})$
$-({\vec{q}}^{2}\mathcal{O}_{1}+\mathcal{O}^{a}_{1}+\mathcal{O}^{b}_{1}-\mathcal{O}^{a}_{3})J^{B}_{21}$
$-[2d{\vec{q}}^{2}\mathcal{O}_{1}+(d+2)\mathcal{O}^{a}_{1}+(d+2)\mathcal{O}^{b}_{1}-\mathcal{O}^{a}_{2}$
$-\mathcal{O}^{b}_{2}-\mathcal{O}^{b}_{3}](J^{B}_{31}+J^{B}_{42})+\mathcal{O}^{b}_{3}J^{B}_{31}+\mathcal{O}^{a}_{3}J^{B}_{42}\},$ (93)
$\mathcal{M}^{(2)}_{c3.8}$= $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{(-d\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{B}_{41}$
$+({\vec{q}}^{2}\mathcal{O}^{b}_{1}+\mathcal{O}^{c}_{3})(J^{B}_{22}+2J^{B}_{32}+J^{B}_{43})$
$+(\mathcal{O}^{b}_{2}-\mathcal{O}^{b}_{1})J^{B}_{21}+[{\vec{q}}^{2}\mathcal{O}_{1}+\mathcal{O}^{a}_{1}-(d+2)\mathcal{O}^{b}_{1}$
$+\mathcal{O}^{b}_{2}+\mathcal{O}^{a}_{3}+\mathcal{O}^{b}_{3}](J^{B}_{31}+J^{B}_{42})+\mathcal{O}^{b}_{2}J^{B}_{31}+\mathcal{O}^{a}_{2}J^{B}_{42}\},$ (94)
$\mathcal{M}^{(2)}_{c3.9}$= $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{(-d\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{B}_{41}$
$+({\vec{q}}^{2}\mathcal{O}^{a}_{1}+\mathcal{O}^{c}_{3})(J^{B}_{22}+2J^{B}_{32}+J^{B}_{43})$
$+(\mathcal{O}^{a}_{2}-\mathcal{O}^{a}_{1})J^{B}_{21}+[{\vec{q}}^{2}\mathcal{O}_{1}-(d+2)\mathcal{O}^{a}_{1}$
$+\mathcal{O}^{b}_{1}+\mathcal{O}^{a}_{2}+\mathcal{O}^{a}_{3}+\mathcal{O}^{b}_{3}](J^{B}_{31}+J^{B}_{42})$
$+\mathcal{O}^{a}_{2}J^{B}_{31}+\mathcal{O}^{b}_{2}J^{B}_{42}\},$ (95)
$\mathcal{M}^{(2)}_{c3.10}$= $~5\mathcal{M}^{(2)}_{c3.6}/.\{J^{B}\to J^{R}\},$
$\mathcal{M}^{(2)}_{c3.11}$= $~5\mathcal{M}^{(2)}_{c3.7}/.\{J^{B}\to J^{R}\},$ (96)
$\mathcal{M}^{(2)}_{c2.12}$= $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{(-d\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{R}_{41}$
$+({\vec{q}}^{2}\mathcal{O}^{b}_{1}+\mathcal{O}^{c}_{3})(J^{R}_{22}+2J^{R}_{32}+J^{R}_{43})$
$+(\mathcal{O}^{b}_{2}-\mathcal{O}^{b}_{1})J^{R}_{21}+[{\vec{q}}^{2}\mathcal{O}_{1}+\mathcal{O}^{a}_{1}-(d+2)\mathcal{O}^{b}_{1}$
$+\mathcal{O}^{a}_{2}+\mathcal{O}^{b}_{2}+\mathcal{O}^{b}_{3}](J^{R}_{31}+J^{R}_{42})$
$+\mathcal{O}^{b}_{3}J^{R}_{31}+\mathcal{O}^{a}_{3}J^{R}_{42}\},$ (97)
$\mathcal{M}^{(2)}_{c3.13}$= $~4\frac{g_{1}^{2}g_{2}^{2}}{f^{4}}(d-3)A_{1}\{(-d\mathcal{O}_{1}+\mathcal{O}_{2}+\mathcal{O}_{3})J^{R}_{41}$
$+({\vec{q}}^{2}\mathcal{O}^{b}_{1}+\mathcal{O}^{c}_{3})(J^{R}_{22}+2J^{R}_{32}+J^{R}_{43})$
$+(\mathcal{O}^{b}_{2}-\mathcal{O}^{b}_{1})J^{R}_{21}+[{\vec{q}}^{2}\mathcal{O}_{1}+\mathcal{O}^{a}_{1}-(d+2)\mathcal{O}^{b}_{1}$
$+\mathcal{O}^{a}_{2}+\mathcal{O}^{b}_{2}+\mathcal{O}^{a}_{3}](J^{R}_{31}+J^{R}_{42})$
$+\mathcal{O}^{a}_{3}J^{R}_{31}+\mathcal{O}^{b}_{3}J^{R}_{42}\}.$ (98)
@@
The coefficients in the above NLO OPE and TPE amplitudes are listed in Tables 7 and 8, respectively. To further evaluate the calculated amplitudes, we should deal with the terms in (54) according to the partial wave we considered. In the $S$-wave, for the terms like $(q\cdot\epsilon_{i})(q\cdot\epsilon_{j}^{*})(\epsilon_{k}\cdot\epsilon_{l}^{*})$, we can make the following substitutions:
@@
Table 5: The values of the products of polarization vectors in the $S$-wave effective potentials with total angular momentum $J=0$, 11, and $2$, respectively Wang:2018atz .
@@
   $\mathcal{O}_{1}$    11    11    11
   $\mathcal{O}_{2}$    11    1-1    11
   $\mathcal{O}_{3}$    33    0    0
Table 6: The coefficients appearing in the contact term amplitudes at NLO [Eqs. (57)-(74)].
@@
$A_{a3.1}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{1}{4}(3D_{b}-E_{b})$ $\frac{1}{4}(3D_{b}-E_{b}$) $\frac{3}{4}(D_{a}+E_{a})$ $\frac{3}{4}(D_{b}+E_{b})$ $\frac{3}{4}(D_{b}+E_{b})$ 0 0
$A_{a3.2}$ $\frac{1}{4}(3D_{a}-E_{a})$ 0 0 $\frac{3}{4}(D_{a}+E_{a})$ 0 0 $\delta_{2}$ $\delta_{2}$
$A_{a3.3(4)}$ $\frac{1}{4}(3D_{b}-E_{b})$ $\frac{1}{4}(3D_{b}-E_{b})$ 0 $\frac{3}{4}(D_{b}+E_{b})$ $\frac{3}{4}(D_{b}+E_{b})$ 0 $0(\delta_{2})$ $\delta_{2}(0)$
$A_{a3.5}$ 0 0 0 0 0 0 0 0
$A_{a3.6}$ 0 0 0 0 0 0 $\delta_{1}$ $\delta_{1}$
$A_{a3.7(8)}$ 0 0 0 0 0 0 $0(\delta_{1})$ $\delta_{1}(0)$
$A_{a3.9}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{a}+E_{a})$
$\frac{1}{4}[(d-3)(D_{b}+E_{b})$
$(D_{a}+E_{a})]$
$\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{a}-3E_{a})$
$\frac{-3}{4}[(d-3)(D_{b}-3E_{b})$
$(D_{a}-3E_{a})]$
0 0
$A_{a3.10}$ $\frac{1}{4}(D_{b}+E_{b})$ 0 0 $\frac{-3}{4}(D_{b}-3E_{b})$ 0 0 $\delta_{1}$ $\delta_{2}$
$A_{a3.11(12)}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{b}+E_{b})$ 0 $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 $0(\delta_{1})$ $\delta_{2}(0)$
$A_{a3.13}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{a}+E_{a})$
$\frac{1}{4}[(d-3)(D_{b}+E_{b})$
$(D_{a}+E_{a})]$
$\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{a}-3E_{a})$
$\frac{-3}{4}[(d-3)(D_{b}-3E_{b})$
$(D_{a}-3E_{a})]$
0 0
$A_{a3.14}$ $\frac{1}{4}(D_{b}+E_{b})$ 0 0 $\frac{-3}{4}(D_{b}-3E_{b})$ 0 0 $\delta_{1}$ $\delta_{2}$
$A_{a3.15(16)}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{b}+E_{b})$ 0 $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 $0(\delta_{1})$ $\delta_{2}(0)$
$A_{a3.(17+18)}$ $D_{a}+E_{a}$ $D_{b}+E_{b}$ $D_{b}+E_{b}$ $D_{a}-3E_{a}$ $D_{b}-3E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{2}$
$A_{a3.(19+20)}$ $D_{a}+E_{a}$ $D_{b}+E_{b}$ $D_{b}+E_{b}$ $D_{a}-3E_{a}$ $D_{b}-3E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{1}$
Table 7: The constants AA (as well as $\omega_{1,2}$) appearing in the OPE amplitudes [Eqs. (75)-(86)].
@@
$I=1$ $\frac{1}{16}$ $\frac{1}{16}$ $\frac{1}{16}$ $\frac{1}{16}$ $\frac{1}{12}$ $\frac{1}{12}$ $\frac{1}{4}$ $\frac{1}{4}$ $\frac{1}{4}$ 0 0
$I=0$ $\frac{3}{16}$ $\frac{3}{16}$ $\frac{3}{16}$ $\frac{3}{16}$ $\frac{1}{4}$ $\frac{1}{4}$ $\frac{3}{4}$ $\frac{3}{4}$ $\frac{3}{4}$ 0 0
$\omega_{1}$ 0 $\delta_{2}(0)$ 0 $(\delta_{1}) 0 0$ 0 0 0 0 0
$\omega_{2}$ 0 $0(\delta_{2})$ 0 $\delta_{1}(0)$ 0 0 0 $\delta_{1}$ $\delta_{2}$ 0 0
Table 8: The constants appearing in the TPE amplitudes [Eqs. (87)-(98)]. Note that we have $A_{51}=A_{15}$.
@@
   $A_{c3.1}$    $\frac{1}{16}$    $\frac{1}{16}$    $\frac{1}{16}$    $\frac{3}{16}$    $\frac{3}{16}$    $\frac{3}{16}$    0    0
   $A_{c3.2}$    $\frac{i}{8}$    $\frac{i}{8}$    0    $\frac{3i}{8}$    $\frac{3i}{8}$    0    $\delta_{2}$    0
   $A_{c3.3}$    $\frac{i}{8}$    $\frac{i}{8}$    0    $\frac{3i}{8}$    $\frac{3i}{8}$    0    0    0
   $A_{c3.4}$    $\frac{i}{8}$    $\frac{i}{8}$    0    $\frac{3i}{8}$    $\frac{3i}{8}$    0    $\delta_{1}$    0
   $A_{c3.5}$    $\frac{i}{8}$    $\frac{i}{8}$    0    $\frac{3i}{8}$    $\frac{3i}{8}$    0    0    0
   $A_{c3.6}$    $\frac{1}{16}$    0    0    $\frac{9}{16}$    0    0    $\delta_{1}$    $\delta_{2}$
   $A_{c3.7}$    $\frac{1}{16}$    0    0    $\frac{9}{16}$    0    0    0    0
   $A_{c3.8}$    $\frac{1}{16}$    0    0    $\frac{9}{16}$    0    0    $\delta_{1}$    0
   $A_{c3.9}$    $\frac{1}{16}$    0    0    $\frac{9}{16}$    0    0    0    $\delta_{2}$
   $A_{c3.10}$    $\frac{5}{16}$    0    0    $\frac{3}{16}$    0    0    $\delta_{1}$    $\delta_{2}$
   $A_{c3.11}$    $\frac{5}{16}$    0    0    $\frac{3}{16}$    0    0    0    0
   $A_{c3.12}$    $\frac{5}{16}$    0    0    $\frac{3}{16}$    0    0    $\delta_{1}$    0
   $A_{c3.13}$    $\frac{5}{16}$    0    0    $\frac{3}{16}$    0    0    0    $\delta_{2}$